OpenStudy (anonymous):

x+5y+1 & 2x+10y=2 solve by elimination need help

OpenStudy (anonymous):

I need help I don't understand this one :(

OpenStudy (gorv):

x+5y=1 x=1-5y 2x+10y=2 put in this x=1-5y

OpenStudy (gorv):

both are same equation

OpenStudy (gorv):

2*(x+5y)=2 x+5y=2/2 x+5y=1

OpenStudy (anonymous):

Y>x+2 Y less than -x+1??